OpenStudy (anonymous):

please help me solve this equation ! 5+4=1/x

OpenStudy (anonymous):

(5+4)x=1 9x=1 x=1/9

OpenStudy (anonymous):

@Jshakour u understood

OpenStudy (anonymous):

yeah kinda, u sure that's right?

OpenStudy (anonymous):

s

OpenStudy (unklerhaukus):

to check\[5+4=\frac{1}{1/9} \]\[9=9\] T

OpenStudy (anonymous):

okay thanks !

OpenStudy (anonymous):

Did you understand the first step? shakir multiplied both sides of the equation by x, which keeps the equality. 5+4 = 1/x (5+4)x = (1/x)x, so (5+4)x=1 Then always check your answer as UnkleRhaukus did

OpenStudy (anonymous):

yeah i understand , thanks !